OpenStudy (anonymous):

How many different 12-member juries can be chosen from a pool of 40 people? I really can't understand this.

sam (.sam.):

40C12?

mathslover (mathslover):

And : \(\large ^n C_r = \cfrac{n!}{r! (n-r)!} \)

mathslover (mathslover):

@Haruhi , can you solve it now?

OpenStudy (anonymous):

not really. I've been trying to solve it that way all this time. I'm getting weird results.

sam (.sam.):

From \[\large ^n C_r = \cfrac{n!}{r! (n-r)!} = \frac{40!}{12!(40-12)!} = \frac{40 \times 39 \times 38 \times...}{(12 \times 11 \times 10 \times ...)(28 \times 27 \times 26 \times ...)}\]

sam (.sam.):

Or just use a calculator

sam (.sam.):

What do you get?

OpenStudy (anonymous):

I get 1.5186643e+48 (whatever that means) and my options are a) 3,586,853,480 b) 4,586,853,480 c) 5,586,853,480 d) 6,586,853,480

sam (.sam.):

Lol you should get 5,586,853,480 check your calculator

OpenStudy (zarkon):

type "calculate 40 choose 12" into google

sam (.sam.):

Ahh

OpenStudy (zarkon):

it will take "40 choose 12" and give the answer

OpenStudy (anonymous):

I think we don't have Google in the exam!?

OpenStudy (anonymous):

Ok..... still confused. How can a calculator be wrong lol what the hell am I doing

OpenStudy (anonymous):

Try:\[\frac{40!}{(12!(40-12)!)} \]

OpenStudy (anonymous):

That is wrong all of you I think its 12 times 40 its simple probability i think

OpenStudy (anonymous):

No nevermind

OpenStudy (anonymous):

Its c

OpenStudy (anonymous):

It's the right answer, yeah. But it's damn confusing having to work with numbers like 815,915,279,999,999,744,392,488,672,368,848,648,400,776,424,296 -_- so I must've made a mistake at some point.

OpenStudy (anonymous):

please dont say dam on a post there are 7th graders in geometry and algebra 2 like me who don't like that

OpenStudy (anonymous):

ok... >_>

OpenStudy (anonymous):

What is that supposed to mean

OpenStudy (anonymous):

weirdos